# Frequency dependence of nonsequential double ionization of atoms in strong laser fields

### v1

@@ I Introduction
Many intriguing and fascinating phenomena in nature thrive from correlations. One of the best example from strong-field physics is so-called nonsequential double or multiple ionization (NSDI or NSMI) of atoms (see e.g. [1, 2] for a review). NSDI and NSMI processes are well described by a three-step model [3] which is more commonly referred to as the rescattering mechanism. The whole story starts with electrons tunneling through the Stark barrier when the laser field is strong. Most of them will give rise to a single ionization signal. A small part of them, however, is accelerated in the laser field and driven back to the ion as the field reverses sign. Finally, in inelastic collisions, the returning electrons transfer some energy to the electrons remaining at the ion. As a result, the electrons form a transient, short-lived compound state which may decay through single, double or multiple ionization events or undergo a repetition of the rescattering cycle [4, 5, 6]. Additionally, the anti-correlated emission caused by a slingshot mechanism is to be expected [7].
In the following, we restrict ourselves to double ionization. There are two main double ionization scenarios [6]. First, there are simultaneously escaping electrons. This channel is dominated by recollision-induced direct ionization (REDI) [8] events. Minor contributions to this channel come from the decays of autoionizing two-electron states that persisted through several field cycles and two-electron tunnel ionizations. Second, there are electrons following a consecutive path of escape, i.e. a path on which there is a time delay between the electrons’ departures. In this channel, the independent escape of electrons, called sequential double ionization, is included. Another important contribution comes from indirect processes where the residual electron is excited to a higher bound state before leaving the atom (recollision-induced excitation plus subsequent field ionization (RESI) [9]). One may also consider RESI-like processes in which the excited state persisted through several field cycles before ionization.
@@
Many theoretical and experimental results concerning NSDI were published within the last few decades addressing different aspects of the phenomenon. Among them is a frequency dependence of NSDI, especially when experimental parameters set the multi-photon ionization regime where the three-step model fails to be valid [21]. On the experimental side, evidence for a frequency dependence of NSDI was already found early [22, 14]. Later, the ratio of double to single ionization was studied for xenon over a broad range of frequencies and intensities [23, 24, 25]. Furthermore, the influence of the frequency on the longitudinal ion momentum distributions of argon and neon was investigated [26, 21]. On the theoretical side, the frequency dependence of NSDI has been studied in several semiclassical calculations [14, 27, 28, 29] and in a quantum mechanical one [30] using a mean-field model [31]. Recently, the frequency scaling of the NSDI yield was studied in joint experimental and semi-classical studies in Xe [32] and Mg [33]. Also the changes in the two-electron momentum distributions were addressed lately in joint experimental and theoretical studies, either from the point of view of $U_{p}$ dependence (the ponderomotive potential, $U_{p}=F_{0}^{2}/4\omega^{2}$[34] or of pulse length dependence [35].
@@ IV Results
Mainly four frequencies will be considered in the following: $0.094$ a.u. (blue), $0.060$ a.u., $0.035$ a.u. and $0.023$ a.u. (all near-infrared), corresponding to wavelengths of 485 nm, 760 nm, 1313 nm and 2000 nm, respectively. Except for the second one, these are just the wavelengths studied in [21]. It is worth noting that the dipole approximation, i.e. assuming that the electric field is independent of position, is well justified in all cases since the wavelengths are much greater than the Bohr radius. The peak field amplitude $F_{0}$ is varied between $0.11$ a.u. and $0.57$ a.u., corresponding to intensities between $4.25\cdot 10^{14}$ W/cm2 and $1.14\cdot 10^{16}$ W/cm2.
@@ IV.1.1 Frequency dependence
The single and double ionization yields (SI and DI, respectively) obtained for $n_{c}=5$, $\varphi=0$ and the four frequencies mentioned above are shown in Fig. 1(a) as a function of the peak field amplitude. Let us first discuss them for a fixed frequency.
@@
[…]
@@
The ratio of double to single ionization is shown in Fig. 1(b), in agreement with experimental data [32, 33]. In the knee regime, the dependence on the field amplitude is weak, giving rise to a plateau-like structure. For $\omega=0.060$ a.u., the ratio is practically constant, an observation also made in [12, 14].
@@
In Fig. 2, we compare the CE and SE yields. For the sake of clarity, the yields belonging to different $\omega$ are artificially separated by a factor of 500. Nevertheless, we observe that they all are of the same order of magnitude below saturation, with the difference becoming more pronounced as $\omega$ grows. For the three lowest frequencies, CE slightly dominates over SE below saturation. This is mainly a consequence of the size and shape of the cruciform region in the configuration space that represents the neutral atom [6]. By adjusting its borders, we can directly influence the magnitude of each channel, as shortly discussed in Section III. The SE yields increase by about three orders of magnitude between 0.11 a.u. and 0.20 a.u. and eventually saturate. In the saturation regime, the slope of the yield (and hence the yield itself) decreases with decreasing $\omega$. The dependence of the CE yields on $\omega$ is even weaker. However, in contrast to the SE case, the CE yields display a knee structure since they contain both RESI and sequential DI, and the later dominates the DI yields above $F_{max}$. It should be noted that there are still enough neutral atoms to allow for NSDI, so the simultaneous yields continue to grow above the knee. Below saturation, the contributions of sequential DI are negligible and the yield in the consecutive channel result mainly from RESI-like processes.
@@ IV.1.2 Influence of the pulse duration
The ionization yields for $\omega=0.060$ a.u., $\varphi=0$ and $n_{c}\in\{1,2,5,10\}$ are shown in Fig. 3(a). Obviously, NSDI is enhanced in a pulse with more field cycles since the height of the knee increases. Such a trend can be easily explained with the three-step model. Since the number of field extrema grows with the number of field cycles, the number of ionized electron wave packets which can take part in rescattering events leading to NSDI increases as well. These wave packets can also move and interfere with each other for a longer time with increasing $n_{c}$ which manifests in the distortion of the yields. This distortion was also visible in the time-dependent probability fluxes studied in [37] where an eight-cycle pulse with a trapezoidal envelope was used. For better illustration of the NSDI enhancement, the ratio of double to single ionization is shown in Fig. 3(b). We find that it increases with increasing pulse duration, at least before sequential double ionization starts to dominate. Moreover, the plateau region grows wider. These trends are in agreement with the experiment in [56] where the ratio was measured for neon and a Ti:sapphire laser ($\omega=0.057$ a.u.) with pulse durations of $12$ fs and $50$ fs.
For case (ii), $F_{sat}$ and $F_{max}$ are given in Table 2. Clearly, they both decrease with growing $n_{c}$. This trend was also found in [30] and can be explained as follows. As apparent from Eq. 4, the saturation field amplitude $F_{sat}$ depends on the frequency and the number of field cycles. In the multiphoton case [55], the static-field ionization rate $W(F_{0})$ increases monotonically with $F_{0}$. Provided that $F_{0}$ is not too big, this also applies to the tunneling case [57]. Let $F(t)=F_{0}\Theta(T_{p}-t)$. Then Eq. 4 reduces to $W(F_{sat})T_{p}=1$ and $F_{sat}$ decreases if the pulse duration is increased. Now let $F(t)=F_{0}\sin(\omega t)\Theta(T_{p}-t)$. In this case, $\omega$ and $n_{c}$ can be scaled out of the integral and Eq. 4 can be written as $(T_{p}/2\pi)\int_{0}^{2\pi}W(F_{sat}\sin(\tau))\mathrm{d}\tau=1$. Again, $F_{sat}$ decreases with increasing $T_{p}$. Finally, let $F(t)$ be defined by Eq. 2. Taking $x=(t-T_{p}/2)/T_{p}$ leads to $T_{p}\int_{-1/2}^{1/2}W(F_{sat}f(x))\mathrm{d}x=1$. Hence, $F_{sat}$ decreases with increasing $T_{p}$ – the longer the pulse, the smaller the field amplitude where the ionization probability reaches unity and the knee regime begins.
@@ IV.1.3 Rate equations
Figure 5: Comparison of the numerical ionization yields to the ones obtained by integrating the rate equations (5) for $n_{c}=5$, $\omega=0.060$ a.u. and $\varphi=0$. The sequential ionization rates ($W_{01}$ and $W_{12}$) are approximated with the ADK formula, the nonsequential one is approximated by $W_{02}=0.019\cdot W_{01}$ (solid lines) or set to zero (dashed lines).
@@
Some of the observations from the previous subsections can be already understood without the nonsequential contribution, i.e. for $W_{02}=0$. Figs. 6(a)–(c) show the purely sequential SI and DI yields obtained by solving the rate equations (5) for the cases (i)–(iii), respectively. The SI maximum exhibits the same behavior as in Figs. 1(a), 3(a) and 4, i.e. it only depends on the pulse duration. As $F_{max}$ marks the regime where sequential DI starts to dominate the DI channel, the DI yields obtained from the rate equations should resemble its behavior. Indeed, the changes in the DI yields with frequency and number of cycles, Figs. 6(a) and (b), respectively, agree very well with the previously discussed shifts in position of $F_{max}$. This observation is pin-pointed by the unchanged DI yields for constant $\omega/n_{c}$, see Fig. 6(c).
@@ IV.2 Momentum distributions
To make our study comprehensive, let us discuss the two-electron momentum distributions, shown in Figs. 710. For each frequency, we picked four field amplitudes to address the different regimes of double ionization visible in Fig. 1(a): $F_{0}=0.16$ a.u. (nonsequential regime, below the knee), $F_{0}=0.20$ a.u., $F_{0}=0.30$ a.u. (saturation regime, the knee) and $F_{0}=0.40$ a.u. (sequential regime, above the knee).
As can be inferred from the Figs. 710, the width of the distributions grows with increasing $F_{0}$ and decreasing $\omega$. This observation is related to the increase of the ponderomotive energy $U_{p}=F_{0}^{2}/4\omega^{2}$ and in line with the experiments (see e.g. [8, 21, 26, 34]). For comparison, the parallel momentum threshold of ±$\pm 2\sqrt{U_{p}}$ predicted by the rescattering mechanism is also included in the figures. It corresponds to a free electron leaving the ion parallel to the field polarization axis with the maximum energy it can classically acquire in the field ($2U_{p}$). Due to the choice of the laser pulse shape, we obtain zero drift momentum and the boundary set by the threshold is just a circle centered around the origin [9]. In all cases, the two-electron momentum distributions lie well within these circles.
@@
Figure 8: (color online) Two-electron momentum distributions for $n_{c}=5$, $\omega=0.060$ a.u., $\varphi=0$ and the field amplitudes (from upper left to lower right) $F_{0}=0.16$ a.u., $F_{0}=0.20$ a.u., $F_{0}=0.30$ a.u. and $F_{0}=0.40$ a.u. The white circles correspond to the classical kinematical constraint ±$\pm 2\sqrt{U_{p}}$ for the electrons’ momenta.
Figure 9: (color online) Two-electron momentum distributions for $n_{c}=5$, $\omega=0.035$ a.u., $\varphi=0$ and the field amplitudes (from upper left to lower right) $F_{0}=0.16$ a.u., $F_{0}=0.20$ a.u., $F_{0}=0.30$ a.u. and $F_{0}=0.40$ a.u. The white circles correspond to the classical kinematical constraint ±$\pm 2\sqrt{U_{p}}$ for the electrons’ momenta.
Figure 10: (color online) Two-electron momentum distributions for $n_{c}=5$, $\omega=0.023$ a.u., $\varphi=0$ and the field amplitudes (from upper left to lower right) $F_{0}=0.16$ a.u., $F_{0}=0.20$ a.u., $F_{0}=0.30$ a.u. and $F_{0}=0.40$ a.u. The white circles correspond to the classical kinematical constraint ±$\pm 2\sqrt{U_{p}}$ for the electrons’ momenta.
@@
Figure 11: (color online) Longitudinal ion momentum distributions for $n_{c}=5$, $F_{0}=0.16$ a.u. and different frequencies $\omega$, averaged over 20 carrier-envelope phases ($\varphi=j\pi/10$, $0\leq j\leq 19$). All distributions are normalized to the same maximum count.
We compare of our numerical results to experimental ones by inspecting the longitudinal ion momentum distributions. Two groups [21, 26] reported such measurements for doubly-ionzed neon and argon for different frequencies in the nonsequential regime. The key observation of those reports was that the ”valley” at zero parallel ion momentum becomes more pronounced with decreasing frequency. Since the lasers used in these experiments were not phase-stabilized, we have to average our momentum distributions over a uniform distribution of carrier-envelope phases to enable qualitative comparison. We picked a peak field amplitude which is in the nonsequential regime for all frequencies ($F_{0}=0.16$ a.u.) and calculated the distributions for ten phases $\varphi=j\pi/10$ with $0\leq j\leq 9$. Exploiting the fact that $\sin(\omega t+\varphi+\pi)=-\sin(\omega t+\varphi)$, i.e. the two-electron momentum distribution for $\varphi+\pi$ is equal to the one for $\varphi$ reflected on the antidiagonal $p_{1}=-p_{2}$, we can average over the 20 phases $\varphi=j\pi/10$ with $0\leq j\leq 19$. The resulting averaged longitudinal ion momentum distributions are presented in Fig. 11. All distributions are normalized to the same maximum count. In agreement with [21, 26], we find that the depth of the valley strongly depends on $\omega$. However, the deepening of the valley with decreasing $\omega$ is not as monotonic as it appeared to be in [21], due to lowered dimensionality of our model. Nevertheless, deeper valley at lower frequency points to smaller role of RESI in the overall double ionization.

### v2

@@ I Introduction
Many intriguing and fascinating phenomena in nature thrive from correlations. One of the best examples from strong-field physics is so-called nonsequential double or multiple ionization (NSDI or NSMI) of atoms (see e.g. [1, 2] for a review). NSDI and NSMI processes are well described by a three-step model [3] which is more commonly referred to as the rescattering mechanism. The whole story starts with electrons tunneling through the Stark barrier when the laser field is strong. Most of them will give rise to a single ionization signal. A small part of them, however, is accelerated in the laser field and driven back to the ion as the field reverses sign. Finally, in inelastic collisions, the returning electrons transfer some energy to the electrons remaining at the ion. As a result, the electrons form a transient, short-lived compound state which may decay through single, double or multiple ionization events or undergo a repetition of the rescattering cycle [4, 5, 6]. Additionally, the anti-correlated emission caused by a slingshot mechanism is to be expected [7].
In the following, we restrict ourselves to double ionization. There are two main double ionization channels [6]. First, there are simultaneously escaping electrons. This channel is dominated by recollision-induced direct ionization (REDI) [8] events. Minor contributions to this channel come from the decays of autoionizing two-electron states that persisted through several field cycles and two-electron tunnel ionizations. Second, there are electrons following a consecutive path of escape, i.e. a path on which there is a time delay between the electrons’ departures. In this channel, the independent escape of electrons, called sequential double ionization, is included. Another important contribution comes from indirect processes where the residual electron is excited to a higher bound state before leaving the atom (recollision-induced excitation plus subsequent field ionization (RESI) [9]). One may also consider RESI-like processes in which the excited state persisted through several field cycles before ionization.
@@
Many theoretical and experimental results concerning NSDI were published within the last few decades addressing different aspects of the phenomenon. On the experimental side, evidence for a frequency dependence of NSDI was already found early [21, 14]. Later, the ratio of double to single ionization was studied for xenon over a broad range of frequencies and intensities [22, 23, 24]. Furthermore, the influence of the frequency on the longitudinal ion momentum distributions of argon and neon was investigated [25, 26]. On the theoretical side, the frequency dependence of NSDI has been studied in several semiclassical calculations [14, 27, 28, 29] and in a quantum mechanical one [30] using a mean-field model [31]. Recently, the frequency scaling of the NSDI yield was studied in joint experimental and semi-classical studies in Xe [32] and Mg [33]. Also the changes in the two-electron momentum distributions were addressed lately in joint experimental and theoretical studies, covering both the dependence on the ponderomotive energy $U_{p}=F_{0}^{2}/4\omega^{2}$ [34] or the pulse duration [35].
@@ IV Results
Four frequencies will be considered in the following: $0.094$ a.u. (blue), $0.060$ a.u., $0.035$ a.u. and $0.023$ a.u. (all near-infrared), corresponding to wavelengths of 485 nm, 760 nm, 1313 nm and 2000 nm, respectively. Except for the second one, these are just the wavelengths studied in [26]. It is worth noting that the dipole approximation, i.e. assuming that the electric field is independent of position, is well justified in all cases since the wavelengths are much greater than the Bohr radius. The peak field amplitude $F_{0}$ is varied between $0.11$ a.u. and $0.57$ a.u., corresponding to intensities between $4.25\cdot 10^{14}$ W/cm2 and $1.14\cdot 10^{16}$ W/cm2.
@@ IV.1.1 Frequency dependence
The single and double ionization yields (SI and DI, respectively) obtained for $n_{c}=5$, $\varphi=0$ and the four frequencies mentioned above are shown in Fig. 1(a) as a function of the peak field amplitude. We first discuss them for a fixed frequency.
@@
Overall, the frequency dependence of the yields in Fig. 1(a) is weak, as opposed to the strong dependence on the peak field amplitude. The yields are of the same order of magnitude for each of the ionic species and over the whole range of field amplitudes. However, the order with respect to $\omega$ differs. On the one hand, the SI yield increases with increasing $\omega$ for the lowest field amplitudes but this order reverses above $F_{0}\approx 0.15$ a.u. On the other hand, the DI yield increases with increasing $\omega$ for nearly all values of $F_{0}$, with $\omega=0.023$ a.u. being the obvious exception to this trend (at higher field amplitudes, above $F_{0}\approx 0.3$ a.u.). The obtained results are in very good agreement with the data available in the literature [32, 33].
@@
The ratio of double to single ionization is shown in Fig. 1(b) and is in agreement with experimental data [32, 33]. In the knee regime, the dependence on the field amplitude is weak, giving rise to a plateau-like structure. For $\omega=0.060$ a.u., the ratio is practically constant, an observation also made in [12, 14].
@@
In Fig. 2, we compare the CE and SE yields. For the sake of clarity, the yields belonging to different $\omega$ are artificially separated by a factor of 500. Nevertheless, we observe that they all are of the same order of magnitude below saturation, with the difference becoming more pronounced as $\omega$ grows. For the three lowest frequencies, CE slightly dominates over SE below saturation. This is mainly a consequence of the size and shape of the cruciform region in the configuration space that represents the neutral atom [6]. By adjusting its borders, we can directly influence the magnitude of each channel, as shortly discussed in Section III. The SE yields increase by about three orders of magnitude between 0.11 a.u. and 0.20 a.u. and eventually saturate. In the saturation regime, the slope of the yield (and hence the yield itself) decreases with decreasing $\omega$. The dependence of the CE yields on $\omega$ is even weaker. However, in contrast to the SE case, the CE yields display a knee structure since they contain both RESI and sequential DI, and the latter dominates the DI yields above $F_{max}$. It should be noted that there are still enough neutral atoms to allow for NSDI, so the simultaneous yields continue to grow above the knee. Below saturation, the contributions of sequential DI are negligible and the yield in the consecutive channel results mainly from RESI-like processes.
@@ IV.1.2 Influence of the pulse duration
For case (ii), we choose $\omega=0.060$ a.u., $\varphi=0$ and $n_{c}\in\{1,2,5,10\}$, and the relevant ionization yields are shown in Fig. 3(a). Obviously, NSDI is enhanced in a pulse with more field cycles since the height of the knee increases. Such a trend can be easily explained with the three-step model. Since the number of field extrema grows with the number of field cycles, the number of ionized electron wave packets which can take part in rescattering events leading to NSDI increases as well. These wave packets can also move and interfere with each other for a longer time with increasing $n_{c}$ which manifests in the distortion of the yields. This distortion was also visible in the time-dependent probability fluxes studied in [37] where an eight-cycle pulse with a trapezoidal envelope was used. For better illustration of the NSDI enhancement, the ratio of double to single ionization is shown in Fig. 3(b). We find that it increases with increasing pulse duration, at least before sequential double ionization starts to dominate. Moreover, the plateau region grows wider. These trends are in agreement with the experiment in [56] where the ratio was measured for neon and a Ti:sapphire laser ($\omega=0.057$ a.u.) with pulse durations of $12$ fs and $50$ fs.
The $F_{sat}$ and $F_{max}$ obtained in the discussed instance are given in Table 2. Clearly, they both decrease with growing $n_{c}$. This trend was also found in [30] and can be explained as follows. As apparent from Eq. 4, the saturation field amplitude $F_{sat}$ depends on the frequency and the number of field cycles. In the multiphoton case [55], the static-field ionization rate $W(F_{0})$ increases monotonically with $F_{0}$. Provided that $F_{0}$ is not too big, this also applies to the tunneling case [57]. Let $F(t)=F_{0}\Theta(T_{p}-t)$. Then Eq. 4 reduces to $W(F_{sat})T_{p}=1$ and $F_{sat}$ decreases if the pulse duration is increased. Now let $F(t)=F_{0}\sin(\omega t)\Theta(T_{p}-t)$. In this case, $\omega$ and $n_{c}$ can be scaled out of the integral and Eq. 4 can be written as $(T_{p}/2\pi)\int_{0}^{2\pi}W(F_{sat}\sin(\tau))\mathrm{d}\tau=1$. Again, $F_{sat}$ decreases with increasing $T_{p}$. Finally, let $F(t)$ be defined by Eq. 2. Taking $x=(t-T_{p}/2)/T_{p}$ leads to $T_{p}\int_{-1/2}^{1/2}W(F_{sat}f(x))\mathrm{d}x=1$. Hence, $F_{sat}$ decreases with increasing $T_{p}$ – the longer the pulse, the smaller the field amplitude where the ionization probability reaches unity and the knee regime begins.
@@ IV.1.3 Rate equations
Figure 5: Comparison of the numerical ionization yields to the ones obtained by integrating the rate equations (5) for $n_{c}=5$, $\omega=0.060$ a.u. and $\varphi=0$. The sequential ionization rates ($W_{01}$ and $W_{12}$) are approximated by the ADK formula, the nonsequential one is approximated by $W_{02}=0.019\cdot W_{01}$ (solid lines) or set to zero (dashed lines).
@@
Some of the observations from the previous subsections can already be understood without the nonsequential contribution, i.e. for $W_{02}=0$. Figs. 6(a)–(c) show the purely sequential SI and DI yields obtained by solving the rate equations (5) for the cases (i)–(iii) (see sec. IV.1.2), respectively. The SI maximum exhibits the same behavior as in Figs. 1(a), 3(a) and 4, i.e. it only depends on the pulse duration. As $F_{max}$ marks the regime where sequential DI starts to dominate the DI channel, the DI yields obtained from the rate equations should resemble its behavior. Indeed, the changes in the DI yields with frequency and number of cycles, Figs. 6(a) and (b), respectively, agree very well with the previously discussed shifts in position of $F_{max}$. This observation is pinpointed by the unchanged DI yields for constant $\omega/n_{c}$, see Fig. 6(c).
@@ IV.2 Momentum distributions
To make our study comprehensive, we discuss the two-electron momentum distributions, shown in Figs. 710. For each frequency, we picked four field amplitudes to address the different regimes of double ionization visible in Fig. 1(a): $F_{0}=0.16$ a.u. (nonsequential regime, below the knee), $F_{0}=0.20$ a.u., $F_{0}=0.30$ a.u. (saturation regime, the knee) and $F_{0}=0.40$ a.u. (sequential regime, above the knee).
As can be inferred from Figs. 710, the width of the distributions grows with increasing $F_{0}$ and decreasing $\omega$. This observation is related to the increase of the ponderomotive energy $U_{p}=F_{0}^{2}/4\omega^{2}$ and in line with the experiments (see e.g. [8, 26, 25, 34]). For comparison, the parallel momentum threshold of ±$\pm 2\sqrt{U_{p}}$ predicted by the rescattering mechanism is also included in the figures. It corresponds to a free electron leaving the ion parallel to the field polarization axis with the maximum energy it can classically acquire in the field ($2U_{p}$). Due to the choice of the laser pulse shape, we obtain zero drift momentum and the boundary set by the threshold is just a circle centered around the origin [9]. In all cases, the two-electron momentum distributions lie well within these circles.
@@
Figure 8: (color online) Same as Fig. 7, but for $\omega=0.060$ a.u.
Figure 9: (color online) Same as Fig. 7, but for $\omega=0.035$ a.u.
Figure 10: (color online) Same as Fig. 7, but for $\omega=0.023$ a.u.
@@
Figure 11: (color online) Longitudinal ion momentum distributions for $n_{c}=5$, $F_{0}=0.16$ a.u. and different frequencies $\omega$, averaged over 20 carrier-envelope phases ($\varphi=2j\pi/20$). All distributions are normalized to the same maximum count.
We compare of our numerical results to experimental ones by inspecting the longitudinal ion momentum distributions. Two groups [26, 25] reported such measurements for doubly-ionzed neon and argon for different frequencies in the nonsequential regime. The key observation of those reports was that the ”valley” at zero parallel ion momentum becomes more pronounced with decreasing frequency. Since the lasers used in these experiments were not phase-stabilized, we have to average our momentum distributions over a uniform distribution of carrier-envelope phases to enable qualitative comparison. We picked a peak field amplitude which is in the nonsequential regime for all frequencies ($F_{0}=0.16$ a.u.) and calculated the distributions for twenty phases $\varphi=2j\pi/20$. The resulting averaged longitudinal ion momentum distributions are presented in Fig. 11. All distributions are normalized to the same maximum count. In agreement with [26, 25], we find that the depth of the valley strongly depends on $\omega$. However, the deepening of the valley with decreasing $\omega$ is not as monotonic as it appeared to be in [26], due to the reduced dimensionality of our model. Nevertheless,the deeper valley at lower frequencies points to a smaller role of RESI in the overall double ionization process.